OpenStudy (anonymous):

A washer and a dryer cost $907 combined. The washer costs $57 more than the dryer. What is the cost of the dryer?

OpenStudy (beccaboo333):

\[907\div2\]

OpenStudy (anonymous):

Then subtract 57?

OpenStudy (anonymous):

396.5

OpenStudy (beccaboo333):

No, add 57

OpenStudy (anonymous):

But then wouldn't you have to subtract the other 57, so it still equals 907

OpenStudy (beccaboo333):

._.

OpenStudy (beccaboo333):

Work with me man!!!

OpenStudy (anonymous):

Sorry lol

OpenStudy (beccaboo333):

Actually.. >.< Now I'm confused

OpenStudy (beccaboo333):

Correct me sir batman >.>

OpenStudy (anonymous):

We did a problem like this yesterday, I'll start you off, and you finish it. x = washer y = dryer x + y = 901 x = y - 57

OpenStudy (anonymous):

Yeah cuz if you add 57 to one, you need to subtract 57 from the dryer so the washer is 57 more than the dryer so it equals 907

OpenStudy (beccaboo333):

>.< shhh I'm tired.. I tried. xD

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

So is it 396.5?

OpenStudy (beccaboo333):

I think so >.>

OpenStudy (anonymous):

(x+57)+y=907 x=washer y=dryer

OpenStudy (anonymous):

@iambatman? What do u think?

OpenStudy (anonymous):

Idk, try doing it with what I gave you :)

OpenStudy (anonymous):

I'm not going to show everything now, you have to put some effort.

OpenStudy (anonymous):

I did, you divide 907 by 2, then if theres a 57 dollar difference, you have to add 57 to the washer a subtract 57 from the dryer, so youd get 396.5

OpenStudy (anonymous):

It's pretty straight forward from what I gave, just a matter of substitution.

OpenStudy (anonymous):

I'm just straight up confused now.

OpenStudy (anonymous):

y - 57 + y =907 2y = 907 +57 2y = 964 y = 482, dryer will cost $482

OpenStudy (anonymous):

But when you add the washer with a 57 dollar difference you get over 1000 dollars

OpenStudy (anonymous):

482=washer 425=dryer

OpenStudy (anonymous):

they equal 907 and the washer is $57 more than the dryer